OpenStudy (anonymous):

Explain how raising a quotient to a power is similar to raising a product to a power.

OpenStudy (anonymous):

In both cases, you multiply the exponents.

OpenStudy (anonymous):

Thaaank You For The Medal xD

OpenStudy (aum):

\(\large (ab)^n = a^nb^n\) \(\Large (\frac ab)^n = \frac{a^n}{b^n}\)